OpenStudy (anonymous):

How do I integrate sec^2 (t) -3^t dt? I know what to do with the sec^2(t) but what do I do with the the -3^t?

OpenStudy (jamesj):

Note first that 3^t = exp((ln 3)t) Hence d/dt of 3^t = ln 3.exp((ln 3)t) = ln 3 . 3^t So now you can integrate 3^t

OpenStudy (anonymous):

so \[\int\limits_{}^{}3^t = \ln3\] ?

OpenStudy (anonymous):

or 3x over that

OpenStudy (jamesj):

What's the integral of e^(ax)? The integral of 3^x is the integral of exp^((ln 3)x)